OpenStudy (anonymous):

The ratio of the height to the width of a packaging label is 5 to 19. If the height of the label is 2 inches, what is its width?

OpenStudy (anonymous):

If the ratio is 5:19, then for each inch of height, how much width do we get? How much then is it for 2 inches? or you can look at it in terms of fractions if you'd like:\[\frac{ 5 }{19 } = \frac{ 2 }{ x }\] solve for x and that's your answer

OpenStudy (calculusxy):

The width will be 15.2. What you do is setup a proportion and 5/19=4/x. Then cross multiply, so 5x=76 and 76 divided by 5 is 15.2

OpenStudy (calculusxy):

If you like my answer you can give me a medal or be my fan.

OpenStudy (anonymous):

x=7.6